# Geometrical properties of strained and twisted moiré heterostructures

### v1

@@ III.3 Monoclinic lattices
Let us now present the generalized geometric description of moiré patterns arising for any two arbitrary monoclinic lattices. We will mostly follow the formalism of Ref. [kogl2023moire]. The lattices vector of the top and bottom layers can be generally written as
@@
The real space moiré vectors $\mathbf{G}_{i}^{R}$ can be computed analytically, with the result [kogl2023moire]
@@
The angle $\beta_{R}$ between the real-space moiré vectors $\mathbf{G}_{1}^{R}$ and $\mathbf{G}_{2}^{R}$ can be calculated as in Eq. (47). Alternatively, one can compute the angle through the relation [kogl2023moire]
@@
where $\hat{\mathbf{G}}_{i}^{R}=\mathbf{G}_{i}^{R}/\left|\mathbf{G}_{i}^{R}\right|$. These formulas are general and hold for any kind of monoclinic lattices and strain (within the limit of small deformations). The particular hexagonal cases follow by taking $\alpha=60^{\circ}$ with $\delta=0$ (homobilayer, Section III.1) and $\delta\neq 0$ (heterobilayer, Section III.2).
@@ III.4 Special moiré patterns
As stated in the previous sections, the interplay between twist and strain can lead to a plethora of moiré pattern geometries. Some special moiré patterns that may be accomplished deserve further discussion. In this section we will review, in particular, three cases in hexagonal homobilayers: (1) One-dimensional moiré patterns, (2) Square moiré patterns, and (3) Hexagonal moiré patterns.
@@ III.4.1 Quasi-unidimensional patterns
The significant deformation of the moiré geometry under strain can lead to a critical situation in which the moiré vectors become collinear [cosma2014moire, sinner2023strain, kogl2023moire, escudero2024designing]. This result in the formation of quasi-unidimensional moiré channels that have been seen in numerous experiments (see Section IV). In reciprocal space, as the critical limit is approached the moiré Brillouin zone is squeezed, up until it collapses at a certain critical strain value. At such critical point, the moiré vector become collinear, and thus are no longer independent. This occurs when the transformation $\mathbf{T}$, which determines the moiré vectors through $\mathbf{g}_{i}=\mathbf{T}\mathbf{b}_{i}$ [Eq. (46)], is such that [sinner2023strain]
@@
Figure 6: Formation of one-dimensional moiré patterns in twisted and strained bilayer graphene. Panels (a) to (c) correspond to a relative twist $\theta=3^{\circ}$ and uniaxial strain with direction $\phi=0^{\circ}$ and magnitudes $\epsilon=0$ , $\epsilon=0.5\epsilon_{c}$ and $\epsilon=\epsilon_{c}$, respectively, where $\epsilon_{c}$ is the critical strain magnitude given by Eq. (87). Reprinted figure with permission from Ref. [sinner2023strain]. Copyright (2023) by the American Physical Society. Last two panels correspond to: (d) A relative twist $\theta=2^{\circ}$ with biaxial strain $\epsilon_{b}=0.5\%$ and shear strain $\epsilon_{s}=\sqrt{\theta^{2}+\epsilon_{b}^{2}}\sim 3.5\%$ [cf. Eq. (88)]; (e) only shear and biaxial strain with magnitudes $\epsilon_{b}=\epsilon_{s}=1.5\%$ [cf. Eq. (90)]. Adapted under the terms of the $C$ BY license from Ref. [escudero2024designing]. Copyright (2024) by the American Physical Society.
In the particular case of uniaxial heterostrain, the collapse condition $\mathrm{det}\,\mathbf{T}=0$ gives a critical strain magnitude [cosma2014moire, sinner2023strain]
@@
independent of the strain angle $\phi$. At low twist angles the critical strain reduces to $\epsilon_{c}\sim\pm\theta/\sqrt{\nu}$ (for graphene $\epsilon_{c}\sim\pm 5\theta/2$). Thus the required critical strain magnitude is relatively small at low twist angles, well within the experimental range. Equation (87) describes with impressive accuracy the formation of unidimensional channels reported in numerous experiments [mendoza2021strain, bai2020excitons].
The critical condition can also take place under other types of strain configurations. For instance, by the correspondence given by Eq. (69) one readily obtains that unidimensional channels can also arise under shear and biaxial strain if [escudero2024designing]
@@
This critical shear strain is $\sim\sqrt{\nu}\approx 0.4$ times smaller than the one required for uniaxial heterostrain. Equation (88) also predicts the possibility of unidimensional channels arising solely by strain (i.e., $\theta=0$), which occurs when [escudero2024designing]
@@
Insights into the moiré Brillouin zone collapse can be obtained by studying the behavior of the moiré vectors as the critical strain limit is approached. In the particular case of one-dimensional channels arising from uniaxial heterostrain along the direction $\phi=0$, for a strain magnitude $\epsilon=x\epsilon_{c}$ with $0<x<1$, the real space moiré vectors $\mathbf{L}_{i}$ can be recasted as [sinner2023strain]
@@
So far we have considered homobilayer configurations. The generalization to a heterobilayer requires replacing $\mathbf{T}$ by $\mathbf{T}_{h}$, cf. Eq. (71). For the particular case of uniaxial heterostrain in two honeycomb lattices with a lattice mismatch $\sim\epsilon_{b}=\delta$, the critical condition $\mathrm{det}\,\mathbf{T}=0$ yields, to lowest order in $\delta$ and $\theta$, a critical strain magnitude [cosma2014moire]
@@
At low twist angles this reduces to Eq. (87) in the homobilayer case ($\delta=0$). As before, this critical strain is independent of the strain angle $\phi$. The above critical strain becomes relevant in moiré heterostructures of, e.g., graphene on top of hBN [yankowitz2012emergence, xue2011scanning, decker2011local], where $\delta\sim 0.016$ gives the lattice mismatch between both layers [cea2020band]. This lattice mismatch actually allows Eq. (99) to be satisfied even without a twist, which implies a minimum strain magnitude $\left|\epsilon_{c}\right|\sim\delta$ [cosma2014moire], see Figure 4(a).
Figure 7: Three ways to generate hexagonal moiré patterns in a hexagonal homobilayer configuration: (a) only twist angle, (b) only shear strain and (c) only biaxial strain. The three examples correspond to a twist $\theta=5^{\circ}$, shear strain $\epsilon_{s}=2\sin\left(\theta/2\right)\approx 8.7\%$ and biaxial strain $\epsilon_{b}=2\sin\left(\theta/2\right)\approx 8.7\%$. By Eq. (103), the three cases have the same moiré periodicity. Below each moiré pattern it is shown a zoom of the twisted or strained hexagonal layers within the moiré unit cell, highlighting the markedly differences in the atomic position of each case, despite their similarities at the moiré scale. Panel (d) schematically shows the twist or strain effect in the top and bottom layers (twist, shear and biaxial strain, from top to bottom); the effects are exaggerated for better visualization. Adapted under the terms of the $C$ BY license from Ref. [escudero2024designing]. Copyright (2024) by the American Physical Society.
@@ III.4.2 Square patterns
Another interesting case is the formation of square patterns. These have been predicted theoretically [kogl2023moire, escudero2024designing] (see Figures 3 and 5), and observed experimentally [mendoza2021strain, tilak2022moire, yu2022tunable] (see Section IV). In general there is a family of twist and strain configurations that lead to square moiré patterns (see, e.g., label II in Figure 5). Geometrically, they can be identified by when the moiré vectors have equal length and are perpendicular. Within the formalism of Ref. [escudero2024designing] (Section III.1), the square condition can be stated in term of the $\mathbf{F}$ matrix as [cf. Eq. (46)]
@@
At low twist angle $\tan\left(\theta/2\right)\sim\theta/2$, so $\epsilon_{\mathrm{sq}}\propto\theta$, in agreement with Figure 5 (as noted in Section III.1.1, this linear dependence at low $\theta$ occurs for any angle $\beta$ between the moiré vectors). For $\nu=0.16$, one has $\epsilon_{\mathrm{sq}}\approx 0.94\tan\left(\theta/2\right)$ and $\phi_{\mathrm{sq}}\approx-9.4^{\circ}$. At low twist angles ($\theta<5\text{°}$), the needed strain magnitude is thus relatively small; close to the magic angle ($\theta\sim 1\text{°}$) one has $\epsilon_{\mathrm{sq}}\sim 0.8\%$. A square moiré pattern formed by uniaxial strain is shown in Figure 3. In Ref. [yu2022tunable] a quasi-square moiré pattern is observed in TBG at $\theta=0.38^{\circ}$ with uniaxial heterostrain $\epsilon\sim\left(0.21\pm 0.12\right)\%$, in good agreement with the theoretical prediction $\epsilon_{\mathrm{sq}}\left(\theta=0.38^{\circ}\right)\approx 0.3\%$ [escudero2024designing].
@@ III.4.3 Hexagonal patterns
Hexagonal moiré patterns emerge naturally in the case of only twist without strain (e.g., in twisted bilayer graphene) [andrei2020graphene, andrei2021marvels]. But one can also have hexagonal moiré patterns with twist and strain [escudero2024designing]. The simplest case is, of course, with twist and biaxial strain, as this always preserves the hexagonal shape or the layers (Section III.1.3). For a strain that deforms the layers, one can still have hexagonal patterns if the strain changes the angle between the moiré vector to $\beta=60^{\circ}$, as this leads to the same periodicity as with only a twist (where $\beta=120^{\circ}$). The crucial difference is that the AA stacking regimes are stretched [bi2019designing, escudero2024designing], so the moiré pattern will actually look different (the larger the strain, the more different); see panel Panel (b) of Figure 3. In the particular case of uniaxial strain, the hexagonal moiré solutions are obtained from Eqs. (57) and (58) with $\beta=60^{0}$
@@
Another configuration that yields hexagonal patterns is with only biaxial or shear strain [cazeaux2023relaxation, escudero2024designing]. Indeed, in both cases the transformation $\mathbf{F}$ is reduced to $\mathbf{F}=\epsilon^{2}\mathbb{I}$ [cf. Eqs. (61) and (62)], where $\epsilon$ is the shear or biaxial strain magnitude. As this transformation $\mathbf{F}$ is a spherical tensor, it leads to hexagonal moiré patterns. Therefore, pure biaxial or shear strain offer a pathway to engineer hexagonal moiré patterns solely by strain. In the case of pure biaxial strain, the moiré vectors are simply rescaled by the strain magnitude, $\mathbf{G}_{i}=\epsilon_{b}\mathbf{b}_{i}$, so the moiré unit cell has the same orientation as the layers unit cell. For shear strain, the orientation of the hexagonal pattern depends on the shear angle $\varphi$.
@@
Figure 7 shows three hexagonal moiré patterns, realized by only a twist, only shear strain and only biaxial strain. All cases correspond to the same moiré periodicity. Although at large moiré scales the three moiré patterns look almost indistinguishable, the actual atomic displacements in each case are markedly different. This turns out to have a significant impact on the electronic properties of the system [escudero2024designing]. In particular, flat bands around a magic twist angle (equivalent or not) only emerge in the no-strain configuration.
@@ III.5 Moiré Brillouin zone
Figure 9: Evolution of the moiré Brillouin zone (mBZ) for different angles $\beta$ between equal-length moiré vectors $\mathbf{G}_{i}$ (black arrows). For each angle $\beta$, the correct mBZ defined by the points $\mathbf{Q}_{i}$ of Eq. (105) is shown in blue, while the stretched irregular hexagon defined by the points $\mathbf{q}_{i}$ of Eq. (104) is shown in red. Only in the no strain case of $\beta=120^{\circ}$ do both cells coincide. For nonzero strain, only the correct mBZ reflects the periodicity of the moiré pattern. Adapted under the terms of the $C$ BY license from Ref. [escudero2024designing]. Copyright (2024) by the American Physical Society.
In the the case of only a twist with no strain, the moiré vector have always a $120^{\circ}$ angle between them, and the moiré pattern is triangular. The corresponding mBZ is thus a hexagon, and it is straightforward to show that its six borders can be determined from the moiré vectors by the points [moon2013optical, koshino2018maximally]
@@
In general, the correct mBZ must be constructed for the specific twist and strain (in practice, numerically). The construction is greatly simplified in the particular case of twist and strain that give equal length moiré vectors (see Section III.1.1). In that case, the correct points that determine the mBZ are given by [escudero2024designing]
@@
The mBZ depicts the symmetry of the twisted and strained moiré vectors. In conventional band theory, the borders of the BZ represent special, high-symmetries points because they are located at the middle between two (or more) Bragg planes, at which a periodic potential opens a gap and there is a constructive interference condition (Bragg reflection). In nonstrained moiré heterostructures, the system preserves the symmetries of the underlying hexagonal lattices, and the borders of the mBZ still represent high-symmetry points. However, under strain the system loses most of its symmetries and the borders of the mBZ do no longer represent, in general, special high-symmetries points. In particular, under strain the positions of the two Dirac points are no longer located at the vertices of the mBZ [naumis2017electronic, bi2019designing, escudero2024designing]; the Dirac points are rather located at unfixed points within the mBZ (their exact position depends on the specific twist and strain parameters). Therefore, unlike conventional band theory in symmetric crystal structures, the borders of the twisted and strained mBZ do not necessarily reflect special points of the moiré band structure (e.g., where the Dirac points are located). The correct shape of the mBZ can still be important in interpreting, and correctly explaining, experimental transport measurements [wang2023unusual].

### v2

@@ III.3 Monoclinic lattices
Let us now present the generalized geometric description of moiré patterns arising for any two arbitrary monoclinic lattices. We will mostly follow the formalism of Ref. [60]. The lattices vector of the top and bottom layers can be generally written as
@@
The real space moiré vectors $\mathbf{G}_{i}^{R}$ can be computed analytically, with the result [60]
@@
The angle $\beta_{R}$ between the real-space moiré vectors $\mathbf{G}_{1}^{R}$ and $\mathbf{G}_{2}^{R}$ can be calculated as in Eq. (47). Alternatively, one can compute the angle through the relation [60]
@@
where $\hat{\mathbf{G}}_{i}^{R}=\mathbf{G}_{i}^{R}/\left|\mathbf{G}_{i}^{R}\right|$. These formulas are general and hold for any kind of monoclinic lattices and strain (within the limit of small deformations). The particular hexagonal cases follow by taking $\alpha=60^{\circ}$ with $\delta=0$ (homobilayer, Section III.1) and $\delta\neq 0$ (heterobilayer, Section III.2). The initial assumption that only the top layer is twisted can be relaxed by setting $\psi_{i}\rightarrow\psi_{i}+\theta_{-}$ in Eq. (77), where $\theta_{-}$ is now twist in the bottom layer. The particular symmetric twist configuration $\pm\theta/2$ follows by taking $\psi_{i}\rightarrow\psi_{i}\mp\theta/2$ in Eqs. (74) and (77), and in all subsequent expressions for the moiré vectors.
@@ III.4 Special moiré patterns
The formalism presented in the previous sections implies that the interplay between twist and strain in stacked two-dimensional materials can lead to many different moiré geometries. Crucially, due to the magnifying effect of the moiré, this plethora of moiré geometries can be archived even with very small strain magnitudes, well within the experimental regime (see Section IV and Table 1). Thus, with the right combination of twist and strain it becomes possible to engineer any particular 2D moiré geometry.
In this section, we use the general formalism of the Sections III.A-C to obtain and describe three relevant moiré geometries that can be realized with twist and strain: (1) One-dimensional moiré patterns, (2) Square moiré patterns, and (3) Hexagonal moiré patterns.
@@ III.4.1 Quasi-unidimensional patterns
The significant deformation of the moiré geometry under strain can lead to a critical situation in which the moiré vectors become collinear [27, 107, 60, 34]. This result in the formation of quasi-unidimensional moiré channels that have been seen in numerous experiments (see Section IV). In reciprocal space, as the critical limit is approached the moiré Brillouin zone is squeezed, up until it collapses at a certain critical strain value. At such critical point, the moiré vector become collinear, and thus are no longer independent. This occurs when the transformation $\mathbf{T}$, which determines the moiré vectors through $\mathbf{g}_{i}=\mathbf{T}\mathbf{b}_{i}$ [Eq. (46)], is such that [107]
@@
Figure 6: Formation of one-dimensional moiré patterns in twisted and strained bilayer graphene. Panels (a) to (c) correspond to a relative twist $\theta=3^{\circ}$ and uniaxial strain with direction $\phi=0^{\circ}$ and magnitudes $\epsilon=0$ , $\epsilon=0.5\epsilon_{c}$ and $\epsilon=\epsilon_{c}$, respectively, where $\epsilon_{c}$ is the critical strain magnitude given by Eq. (87). Reprinted figure with permission from Ref. [107]. Copyright (2023) by the American Physical Society. Last two panels correspond to: (d) A relative twist $\theta=2^{\circ}$ with biaxial strain $\epsilon_{b}=0.5\%$ and shear strain $\epsilon_{s}=\sqrt{\theta^{2}+\epsilon_{b}^{2}}\sim 3.5\%$ [cf. Eq. (88)]; (e) only shear and biaxial strain with magnitudes $\epsilon_{b}=\epsilon_{s}=1.5\%$ [cf. Eq. (90)]. Adapted under the terms of the $C$ BY license from Ref. [34]. Copyright (2024) by the American Physical Society.
In the particular case of uniaxial heterostrain, the collapse condition $\mathrm{det}\,\mathbf{T}=0$ gives a critical strain magnitude [27, 107]
@@
independent of the strain angle $\phi$. At low twist angles the critical strain reduces to $\epsilon_{c}\sim\pm\theta/\sqrt{\nu}$ (for graphene $\epsilon_{c}\sim\pm 5\theta/2$). Thus the required critical strain magnitude is relatively small at low twist angles, well within the experimental range. Equation (87) describes with impressive accuracy the formation of unidimensional channels reported in numerous experiments [80, 7].
The critical condition can also take place under other types of strain configurations. For instance, by the correspondence given by Eq. (69) one readily obtains that unidimensional channels can also arise under shear and biaxial strain if [34]
@@
This critical shear strain is $\sim\sqrt{\nu}\approx 0.4$ times smaller than the one required for uniaxial heterostrain. Equation (88) also predicts the possibility of unidimensional channels arising solely by strain (i.e., $\theta=0$), which occurs when [34]
@@
Insights into the moiré Brillouin zone collapse can be obtained by studying the behavior of the moiré vectors as the critical strain limit is approached. In the particular case of one-dimensional channels arising from uniaxial heterostrain along the direction $\phi=0$, for a strain magnitude $\epsilon=x\epsilon_{c}$ with $0<x<1$, the real space moiré vectors $\mathbf{L}_{i}$ can be recasted as [107]
@@
So far we have considered homobilayer configurations. The generalization to a heterobilayer requires replacing $\mathbf{T}$ by $\mathbf{T}_{h}$, cf. Eq. (71). For the particular case of uniaxial heterostrain in two honeycomb lattices with a lattice mismatch $\sim\epsilon_{b}=\delta$, the critical condition $\mathrm{det}\,\mathbf{T}=0$ yields, to lowest order in $\delta$ and $\theta$, a critical strain magnitude [27]
@@
At low twist angles this reduces to Eq. (87) in the homobilayer case ($\delta=0$). As before, this critical strain is independent of the strain angle $\phi$. The above critical strain becomes relevant in moiré heterostructures of, e.g., graphene on top of hBN [126, 124, 31], where $\delta\sim 0.016$ gives the lattice mismatch between both layers [24]. This lattice mismatch actually allows Eq. (99) to be satisfied even without a twist, which implies a minimum strain magnitude $\left|\epsilon_{c}\right|\sim\delta$ [27], see Figure 4(a).
Figure 7: Three ways to generate hexagonal moiré patterns in a hexagonal homobilayer configuration: (a) only twist angle, (b) only shear strain and (c) only biaxial strain. The three examples correspond to a twist $\theta=5^{\circ}$, shear strain $\epsilon_{s}=2\sin\left(\theta/2\right)\approx 8.7\%$ and biaxial strain $\epsilon_{b}=2\sin\left(\theta/2\right)\approx 8.7\%$. By Eq. (103), the three cases have the same moiré periodicity. Below each moiré pattern it is shown a zoom of the twisted or strained hexagonal layers within the moiré unit cell, highlighting the markedly differences in the atomic position of each case, despite their similarities at the moiré scale. Panel (d) schematically shows the twist or strain effect in the top and bottom layers (twist, shear and biaxial strain, from top to bottom); the effects are exaggerated for better visualization. Adapted under the terms of the $C$ BY license from Ref. [34]. Copyright (2024) by the American Physical Society.
@@ III.4.2 Square patterns
Another interesting case is the formation of square patterns. These have been predicted theoretically [60, 34] (see Figures 3 and 5), and observed experimentally [80, 111, 128] (see Section IV). In general there is a family of twist and strain configurations that lead to square moiré patterns (see, e.g., label II in Figure 5). Geometrically, they can be identified by when the moiré vectors have equal length and are perpendicular. Within the formalism of Ref. [34] (Section III.1), the square condition can be stated in term of the $\mathbf{F}$ matrix as [cf. Eq. (46)]
@@
At low twist angle $\tan\left(\theta/2\right)\sim\theta/2$, so $\epsilon_{\mathrm{sq}}\propto\theta$, in agreement with Figure 5 (as noted in Section III.1.1, this linear dependence at low $\theta$ occurs for any angle $\beta$ between the moiré vectors). For $\nu=0.16$, one has $\epsilon_{\mathrm{sq}}\approx 0.94\tan\left(\theta/2\right)$ and $\phi_{\mathrm{sq}}\approx-9.4^{\circ}$. At low twist angles ($\theta<5\text{°}$), the needed strain magnitude is thus relatively small; close to the magic angle ($\theta\sim 1\text{°}$) one has $\epsilon_{\mathrm{sq}}\sim 0.8\%$. A square moiré pattern formed by uniaxial strain is shown in Figure 3. In Ref. [128] a quasi-square moiré pattern is observed in TBG at $\theta=0.38^{\circ}$ with uniaxial heterostrain $\epsilon\sim\left(0.21\pm 0.12\right)\%$, in good agreement with the theoretical prediction $\epsilon_{\mathrm{sq}}\left(\theta=0.38^{\circ}\right)\approx 0.3\%$ [34].
@@ III.4.3 Hexagonal patterns
Hexagonal moiré patterns emerge naturally in the case of only twist without strain (e.g., in twisted bilayer graphene) [4, 3]. But one can also have hexagonal moiré patterns with twist and strain [34]. The simplest case is, of course, with twist and biaxial strain, as this always preserves the hexagonal shape or the layers (Section III.1.3). For a strain that deforms the layers, one can still have hexagonal patterns if the strain changes the angle between the moiré vector to $\beta=60^{\circ}$, as this leads to the same periodicity as with only a twist (where $\beta=120^{\circ}$). The crucial difference is that the AA stacking regimes are stretched [8, 34], so the moiré pattern will actually look different (the larger the strain, the more different); see panel Panel (b) of Figure 3. In the particular case of uniaxial strain, the hexagonal moiré solutions are obtained from Eqs. (57) and (58) with $\beta=60^{0}$
@@
Another configuration that yields hexagonal patterns is with only biaxial or shear strain [22, 34]. Indeed, in both cases the transformation $\mathbf{F}$ is reduced to $\mathbf{F}=\epsilon^{2}\mathbb{I}$ [cf. Eqs. (61) and (62)], where $\epsilon$ is the shear or biaxial strain magnitude. As this transformation $\mathbf{F}$ is a spherical tensor, it leads to hexagonal moiré patterns. Therefore, pure biaxial or shear strain offer a pathway to engineer hexagonal moiré patterns solely by strain. In the case of pure biaxial strain, the moiré vectors are simply rescaled by the strain magnitude, $\mathbf{G}_{i}=\epsilon_{b}\mathbf{b}_{i}$, so the moiré unit cell has the same orientation as the layers unit cell. For shear strain, the orientation of the hexagonal pattern depends on the shear angle $\varphi$.
@@
Figure 7 shows three hexagonal moiré patterns, realized by only a twist, only shear strain and only biaxial strain. All cases correspond to the same moiré periodicity. Although at large moiré scales the three moiré patterns look almost indistinguishable, the actual atomic displacements in each case are markedly different. This turns out to have a significant impact on the electronic properties of the system [34]. In particular, flat bands around a magic twist angle (equivalent or not) only emerge in the no-strain configuration.
@@ III.5 Moiré Brillouin zone
Figure 9: Evolution of the moiré Brillouin zone (mBZ) for different angles $\beta$ between equal-length moiré vectors $\mathbf{G}_{i}$ (black arrows). For each angle $\beta$, the correct mBZ defined by the points $\mathbf{Q}_{i}$ of Eq. (105) is shown in blue, while the stretched irregular hexagon defined by the points $\mathbf{q}_{i}$ of Eq. (104) is shown in red. Only in the no strain case of $\beta=120^{\circ}$ do both cells coincide. For nonzero strain, only the correct mBZ reflects the periodicity of the moiré pattern. Adapted under the terms of the $C$ BY license from Ref. [34]. Copyright (2024) by the American Physical Society.
In the the case of only a twist with no strain, the moiré vector have always a $120^{\circ}$ angle between them, and the moiré pattern is triangular. The corresponding mBZ is thus a hexagon, and it is straightforward to show that its six borders can be determined from the moiré vectors by the points [84, 61]
@@
In general, the correct mBZ must be constructed for the specific twist and strain (in practice, numerically). The construction is greatly simplified in the particular case of twist and strain that give equal length moiré vectors (see Section III.1.1). In that case, the correct points that determine the mBZ are given by [34]
@@
The mBZ depicts the symmetry of the twisted and strained moiré vectors. In conventional band theory, the borders of the BZ represent special, high-symmetries points because they are located at the middle between two (or more) Bragg planes, at which a periodic potential opens a gap and there is a constructive interference condition (Bragg reflection). In nonstrained moiré heterostructures, the system preserves the symmetries of the underlying hexagonal lattices, and the borders of the mBZ still represent high-symmetry points. However, under strain the system loses most of its symmetries and the borders of the mBZ do no longer represent, in general, special high-symmetries points. In particular, under strain the positions of the two Dirac points are no longer located at the vertices of the mBZ [86, 8, 34]; the Dirac points are rather located at unfixed points within the mBZ (their exact position depends on the specific twist and strain parameters). Therefore, unlike conventional band theory in symmetric crystal structures, the borders of the twisted and strained mBZ do not necessarily reflect special points of the moiré band structure (e.g., where the Dirac points are located). The correct shape of the mBZ can still be important in interpreting, and correctly explaining, experimental transport measurements [119].